### v3

@@ Abstract
Leveraging structure from the randomization process, a design-based model of an experiment with a binary intervention and outcome can reveal evidence beyond the average effect without additional data. Our proposed statistical decision rule yields a design-based maximum likelihood estimate (MLE) of the joint distribution of potential outcomes in intervention and control, specified by the numbers of always takers, compliers, defiers, and never takers in the sample. With a visualization, we explain why the likelihood varies with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions. We illustrate how the MLE varies with all possible data in samples of 50 and 200: when the estimated average effect is positive, the MLE includes defiers if takeup is below half in control and above half in intervention, unless takeup is zero in control or full in intervention. Under optimality conditions, for increasing sample sizes in which exhaustive grid search is possible, our rule’s performance increases relative to a rule that places equal probability on all numbers of defiers within the estimated Fréchet bounds. We offer insights into effect heterogeneity in two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Our 95% smallest credible sets for defiers include zero and the estimated upper Fréchet bound, demonstrating that evidence is weak. Yet, our rule yields no defiers in one experiment. In the other, our rule yields the estimated upper Fréchet bound on defiers—a count representing over 18% of the sample.
Keywords: econometrics, statistical decision theory, experiments, health
JEL Codes: C1,C44,C9, I1
@@ 1 Introduction
We begin with a stylized example that demonstrates how a design-based model of an experiment reveals evidence beyond the average effect. In Section 2, we formalize the model and present a likelihood function for the joint distribution of potential outcomes in the sample, which we use to construct a maximum likelihood decision rule. In Section 3, we apply our maximum likelihood decision rule to two real-world experiments. Section 4 discusses our contributions to the literature, which we reserve until the end for clarity. Section 5 presents implications for econometric and applied research.
@@
Suppose you are considering an intervention to increase takeup of a desired health behavior. For example, the desired health behavior could be to quit smoking or to become an organ donor, and the intervention could be a payment or a behavioral nudge. You want to estimate the average effect of the intervention. You would also like to decide if the intervention backfires, and if so, for how many people. Your colleague tells you that an experiment can be informative about the population average effect of the intervention, but it cannot provide evidence that the intervention backfires if the estimated average effect is positive. Instead of considering the population, you focus your attention on effects within the fixed set of subjects in an experiment.
@@
Did the intervention backfire for anyone? Why did you observe the data you observed? What would you have seen under the other possible randomized assignments? In more technical terms, what is the joint distribution of “potential outcomes” (Neyman, 1923; Rubin, 1974, 1977) in the sample? Counterfactual questions like these have attracted recent interest in the study of causal inference (Gelman and Imbens, 2013; Pearl and Mackenzie, 2018; Imbens, 2020; Dawid and Musio, 2022).
@@
We assume as in design-based inference that the potential outcomes of each person are fixed, and therefore the joint distribution of potential outcomes in the sample is fixed, before randomization. Following Angrist, Imbens, and Rubin (1996), we classify people into types based on their potential outcomes in intervention and control: “always takers” take up in intervention and control (effect=0), “compliers” take up in intervention but not control (effect=1), “defiers” do not take up in intervention but take up in control (effect=-1), and “never takers” do not take up in intervention or control (effect=0). We specify the joint distribution of potential outcomes by enumerating the numbers of always takers, compliers, defiers, and never takers in the sample. The joint distribution of potential outcomes provides more information than the average effect; it provides the full distribution of effects and disambiguates between zero effects from always and never takers. What is your best guess of the numbers of always takers, compliers, defiers, and never takers in the experiment you ran with six people?
@@
We decide with our design-based maximum likelihood decision rule that the sample includes four compilers and two defiers, and we develop a novel visualization of potential outcomes to illustrate the logic behind our answer in Figure 1. We represent each of the six people in the sample with a circle. The left half of each circle represents the potential outcome in intervention and the right half represents the potential outcome in control. Orange represents “takeup,” white represents “no takeup,” and grey represents “unobserved.”
Figure 1: In an Experiment with Six People, the Likelihood Varies Among Joint Distributions of Potential Outcomes with the Same Marginal Distributions as the Data (2/3 Takeup in Intervention, 1/3 Takeup in Control) and is Maximized with Four Compilers and Two Defiers
@@
Rectangles with lines down the middle represent experiments with intervention on the left and control on the right. Inside each experiment, assignment to intervention reveals the potential outcome in intervention—the left of the rectangle reveals the left of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in control is unobserved, so it is grey. Similarly, assignment to control reveals the potential outcome in control–the right of the rectangle reveals the right of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in intervention is unobserved, so it is grey.
@@
The data from your experiment, depicted in the box at the top of Figure 1 shows takeup among two of three people in intervention and one of three people in control. Therefore, the estimated marginal distribution of the potential outcome in intervention is 2/3 takeup, and the estimated marginal distribution of the potential outcome in control is 1/3 takeup. These two marginal distributions define the “estimated Fréchet set.” We refer to all joint distributions of potential outcomes with the same marginal distributions as members of the same “Fréchet set.”
@@
The three rows of Figure 1 give the three joint distributions of potential outcomes with the same marginal distributions as the estimated Fréchet set: half circles in which the potential outcomes are unobserved in the data are filled in so that they have the same marginal distributions as the data. Takeup occurs among 2/3 of people in intervention (four of the six circles have orange on the left) and among 1/3 of people in control (two of the six circles have orange on the right). As shown in the three rows, there are three ways to combine the colored half circles into colored full circles, each yielding a different number of defiers (circles with white on the left and orange on the right). This depiction demonstrates how the estimated Fréchet set determines the estimated Fréchet bounds on defiers, which extend here from a lower bound of 0 to an upper bound of 2. How do you decide among these three answers?
@@
The joint distribution of potential outcomes in the first row is consistent with the Imbens and Angrist (1994) assumption of “monotonicity” in the sample, whereby there can be compliers or defiers, but not both. Since the estimated effect of 1/3 is positive, you assume away defiers. You then deduce that the person who takes up in control must be an always taker and the person who does not take up in intervention must be a never taker. Assuming that the randomization guarantees that the count of each type is the same in each arm, the sample includes two always takers and two never takers. The remaining two people must be compliers, so the joint distribution of potential outcomes consistent with the monotonicity assumption includes two always takers, two compliers, and two never takers. However, you are uncomfortable assuming monotonicity because you are concerned that there could be defiers, and you would prefer to decide based on evidence.
@@
The joint distributions of potential outcomes in the next two rows include defiers. The one in the second row includes one always taker, three compliers, one defier, and one never taker. The one in the third row includes four compliers and two defiers. There is a rationale in the literature for deciding on the second row—it has one defier, the midpoint (Li and Pearl, 2019) of the Fréchet bounds on defiers. It can therefore seem surprising that we decide on the distribution in the third row.
@@
The top half of each row in Figure 1 provides a rationale for our decision using a “design-based model” that combines the design of the randomization with a model of potential outcomes to specify a data generating process and use it to generate all possible data. As depicted in the top of each row, there are 20 possible randomized assignments of six people that result in three in intervention ($6\text{ choose }3=20$). Each of the six people in the experiment has a separate row. The person’s type and randomized assignment to intervention or control determines the data the experiment reveals, which can be counterfactual to what you observed. The assignments that yield the data you observed—2/3 takeup in intervention and 1/3 takeup in control—are in lighter shading.
@@
The number of randomized assignments that yield the data you observed gives the numerator of the design-based likelihood, reported in the third column. Paraphrasing the baby board book “Statistical Physics for Babies” (Ferrie, 2017), which provides style inspiration for our visualization, physicists refer to the number of ways that you could have seen what you have seen as “entropy.” The design-based likelihood is equal to the entropy—the number of assignments that yield the data (shown in lighter shading)—divided by the number of possible assignments. The first row, with zero defiers, yields the data in eight of 20 possible assignments, so the likelihood is $8/20=40\%$. The second row, with one defier, yields the data in six of 20 possible assignments, so the likelihood is $6/20=30\%$. The last row, with two defiers, yields the data in 12 of 20 possible assignments, so the likelihood is $12/20=60\%$.
@@
Why do some joint distributions of potential outcomes yield the data in more randomized assignments than others and therefore have higher likelihoods than others? The bottom half of each row sorts the six people by type and separately shows how the always takers, compliers, defiers, and never takers must be randomized into intervention and control to yield the data. In the first row, there are three types—always takers, compliers, and never takers—and there must be a balance between intervention and control within each type to yield the data. In the second row, there are four types—always takers, compliers, defiers, and never takers—and there must be an imbalance between intervention and control within each type to yield the data. In the third row, there are only two types—compliers and defiers—and there must be a balance between intervention and control within each type to yield the data.
@@
People can be the same or different, but the likelihoods are higher when the people who yield the data are as similar as possible in two senses: they belong to fewer types, and those of the same type are balanced between intervention and control. [^1] Andrew Gelman and Keith O-Rourke discuss the importance of “similarity” in statistical evidence: “Awareness of commonness can lead to an increase in evidence regarding the target; disregarding commonness wastes evidence; and mistaken acceptance of commonness destroys otherwise available evidence. It is the tension between these last two processes that drives many of the theoretical and practical controversies within statistics” (Gelman and O’Rourke, 2017). The people who yield the data are most similar in the third row: they belong to the fewest number of types (only two), and those of the same type are balanced between intervention and control. Comparison of the first and third rows illustrates how having fewer types can increase the likelihood. In the first row, there two ways to randomize one of two always takers into intervention (2 $2\text{ choose }1$=2), two ways to randomize one of two compliers into intervention (2 $2\text{ choose }1$=2), and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are eight assignments that yield the data ($2\times 2\times 2=8$). In the third row, there are six ways to randomize two of four compliers into intervention ($4\text{ choose }2=6$) and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are 12 assignments that yield the data ($6\times 2=12$). All assignments that yield the data in the first row also yield the data in the third row, which has fewer types. In the third row, having four compliers makes it possible for four additional assignments—assignments in which the first two people are in the same arm—to yield the data. Joint distributions of potential outcomes other than the ones in Figure 1 can yield the data, but by exhaustive grid search, you determine that the global maximizer of the likelihood is the one depicted in the third row, which only includes two types.
@@
By our proposed maximum likelihood decision rule, you conclude that your best guess for the joint distribution of potential outcomes in the sample consists of four compilers and two defiers. Using this distribution, you recover the estimated average effect of 1/3—of the six people in the experiment, the four compliers have an effect of positive one, and the two defiers have an effect of negative one ($(4\times 1+2\times-1)/6=1/3$). And, while you know the experimental data provides no information about the population share for whom the experiment would backfire, you are able to estimate that, within your sample, the intervention backfired for two subjects. The design-based model of the experiment enables you to learn about not just the average effect of your intervention, but also about something more primitive: the joint distribution of potential outcomes, which yields the entire distribution of effects.
@@ 2 A Design-Based Model of an Experiment
In this section, we formalize the design-based model of an experiment presented in our stylized example. A sample consists of $n$ subjects, each of whom has a binary potential outcome in intervention $Y_{I}\in\{0,1\}$ and a binary potential outcome in control $Y_{C}\in\{0,1\}$, where $1$ represents takeup and 0 represents no takeup. A subject’s observed outcome depends only on their own potential outcomes and their inclusion in the intervention or control arm, ruling out network-type effects through a “no interference” (Cox, 1958) or “stable unit treatment value” (Rubin, 1980) assumption.
@@
where $\mathbf{1}_{\{\bullet\}}$ is the indicator function.
Subjects in the sample belong to one of four types or “principal strata,” defined by combinations of potential outcomes (Frangakis and Rubin, 2002). Let $\theta_{11}$ represent the number of always takers $(Y_{I}=1,Y_{C}=1)$, $\theta_{10}$ the number of compilers $(Y_{I}=1,Y_{C}=0)$, $\theta_{01}$ the number of defiers $(Y_{I}=0,Y_{C}=1)$, and $\theta_{00}$ the number of never takers $(Y_{I}=0,Y_{C}=0)$. The collection of these four counts ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ constitutes the joint distribution of potential outcomes in the sample, which we write as counts rather than shares to ease the following likelihood notation. The experimental data ${\boldsymbol{X}}=(X_{I1},X_{I0},X_{C1},X_{C0})$ consist of the counts of subjects who take up in intervention $X_{I1}$, who do not take up in intervention $X_{I0}$, who take up in control $X_{C1}$, and who do not take up in control $X_{C0}$.
We adopt a “design-based” approach, in which the joint distribution of potential outcomes $\boldsymbol{\theta}$} is fixed, but unknown, and all randomness in the experimental data $\boldsymbol{X}$} comes from the random assignment of subjects to either intervention or control. As shown in Appendix B, randomization implies a distribution for the experimental data $\boldsymbol{X}$} given the joint distribution of potential outcomes $\boldsymbol{\theta}$}, yielding a likelihood expression. For Bernoulli randomization, also known as simple randomization or independent and identically distributed (IID) or randomization, with probability $p$ of assignment to intervention, the likelihood is:
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$ (2)
This last expression appears in Copas (1973).
@@
In the previous section, we present a likelihood for the joint distribution of potential outcomes in the sample. Each joint distribution of potential outcomes implies marginal distributions of potential outcomes in the intervention and control arms. The marginal distribution of the potential outcome in intervention is:
@@
The marginal distribution of the potential outcome in control is:
@@
Given a pair of marginal distributions of potential outcomes in intervention and control, there are numerous joint distributions consistent with them. Boole (1854), Hoeffding (1940), and Fréchet (1957) derive bounds on the possible copulas connecting marginal distributions of random variables into joint distributions. By applying these bounds to a pair of marginal distributions of potential outcomes, we can determine the set of all joint distributions of potential outcomes consistent with the pair of marginal distributions. We refer to the set of joint distributions of potential outcomes consistent with a given pair of marginal distributions as the “Fréchet set” of those marginal distributions, $\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)$. We index the distributions in a Fréchet set by their number of defiers, $\theta_{01}$, which can take values in the range
$\max\{0,-(\theta_{1\bullet}-\theta_{\bullet 1})\}\leq\theta_{01}\leq\min\{\theta_{\bullet 1},n-\theta_{1\bullet}\}$ (3)
These inequalities provide the so-called “Fréchet bounds” on the number of defiers. Within these bounds, each value for the number of defiers implies a unique number of always takers, never takers, and compilers consistent with the given marginals:
@@
$\theta_{00}=n-\theta_{1\bullet}-\theta_{01}\Big\}$
@@
$\hat{\theta}_{1\bullet}({\boldsymbol{x}})$ $=\Big(\frac{n}{\mathbb{P}(\text{Assignment to Intervention})}\Big)\Big(\frac{x_{I1}}{x_{I1}+x_{I0}}\Big)$
$\hat{\theta}_{\bullet 1}({\boldsymbol{x}})$ $=\Big(\frac{n}{\mathbb{P}(\text{Assignment to Control})}\Big)\Big(\frac{x_{C1}}{x_{C1}+x_{C0}}\Big).$ (4)
@@
When we report these estimates, we round them if necessary for comparison to our count-based distributions. In our design-based model of an experiment, these are unbiased estimators of the sample marginal distributions of potential outcomes (before rounding). In a standard sampling-based model of an experiment, these estimators are both unbiased and consistent estimates of the population marginal distributions of potential outcomes.
@@
It is well-known that, in the sampling-based model of an experiment, the takeup counts in intervention and control contain no information about the joint distribution of potential outcomes in the population beyond their marginal distributions. In Appendix C, we recreate this result by showing that the likelihood of the population-level joint distribution of potential outcomes is constant, conditional on the population-level marginal distributions. In other words, the likelihood function is flat across every population-level Fréchet set.
@@
In our design-based model of an experiment, however, the data does contain information about the sample joint distribution of potential outcomes beyond the sample marginal distributions. That is, the likelihood function varies with the number of defiers within sample-level Fréchet sets. Figure 2 plots the likelihood values of each joint distribution of potential outcomes in the Fréchet set formed by the estimated marginal distributions from the second empirical example that we will discuss in Section 3. We index the joint distributions in this Fréchet set by their number of defiers along the horizontal axis. The height of the bar is the likelihood of the joint distribution in the Fréchet set with the specified number of defiers. Within this Fréchet set, the likelihood is highest at the distribution with 21 defiers, the upper Fréchet bound on defiers. As the number of defiers increases, the likelihood initially decreases, and then increases. In all empirical examples we have considered, the likelihood maintains a U-shape within the estimated Fréchet set and all other Fréchet sets, and is maximal at either the minimum or maximum number of defiers. If we normalize likelihoods in this Fréchet set to sum to one, the smallest collection of likelihoods to contain 95% of the mass excludes 8 and 9 defiers, indicated by their lighter shading.

### v4

@@ Abstract
We estimate the numbers of always takers, compliers, defiers, and never takers in the sample of people in an experiment, rather than a hypothetical population from which it was drawn, using structure from the randomization design. Our data include only a binary intervention and outcome. We develop a visualization to show that samples with defiers can sometimes generate the data in more ways than samples without defiers, yielding a higher design-based likelihood. We propose a maximum likelihood decision rule that can harness this evidence, which is not captured by standard hypothesis tests, and we provide optimality conditions. We illustrate the output of our decision rule for all possible data in samples of 50 and 200 with half in intervention, demonstrating a pattern in when the MLE includes defiers despite a positive average effect. We provide insights into effect heterogeneity in two published experiments with interventions that could plausibly backfire for some people despite statistically significant positive average effects on takeup of desirable health behaviors. In both, our 95% credible sets include the estimated Fréchet bounds, demonstrating that evidence is weak. Yet, our MLE includes no defiers in one; in the other, the MLE includes a count of defiers equal to the estimated upper Fréchet bound, over 18% of the sample. The MLE can support a monotonicity assumption or a specific alternative as a step toward improving the average effect of future interventions by targeting them away from noncompliers. Our dbmle package, compatible with Python and Stata, implements our statistics.
Keywords: health, econometrics, statistical decision theory, experiments
JEL Codes: I1, C1, C44, C9
@@ 1 Introduction
We begin with a stylized example that demonstrates how a design-based model of an experiment can reveal evidence beyond the average effect. In Section 2, we formalize the model and present a likelihood function for the joint distribution of potential outcomes in the sample, which we use to construct a maximum likelihood decision rule. In Section 3, we apply our maximum likelihood decision rule to two real-world experiments. Section 4 discusses our contributions to the literature. Section 5 presents implications for econometric and applied research.
@@
Suppose you are considering an intervention to increase takeup of a desired health behavior. For example, the desired health behavior could be to quit smoking, and the intervention could be a payment. You are worried that your intervention could backfire, even if it increases takeup on average. Some people who would otherwise quit smoking could be dissuaded from doing so by a payment. You are also concerned that you would rather not waste payments on people who would quit smoking or continue smoking regardless.
A colleague tells you that you can use an experiment to estimate the average effect of the intervention in the population, but it cannot provide evidence that the intervention backfires if the average effect is positive. You are worried that even if the estimated average effect is positive, the intervention could backfire for some people and be ineffectual for others. Rather than focusing on the average effect in the population, you focus your attention on the distribution of effects within the fixed sample of subjects in an experiment.
@@
Why did you observe the data you observed? What would you have seen under the other possible randomized assignments? Can the experiment reveal any evidence about effect heterogeneity? In more technical terms, what is the joint distribution of “potential outcomes” (neyman1923; rubin1974; rubin1977) in the sample? Counterfactual questions like these have attracted recent interest in the study of causal inference (gelman2013; pearl2018; imbens2020; dawid2022).
@@
We assume, as in design-based inference, that the potential outcomes of each person are fixed, and therefore the joint distribution of potential outcomes in the sample is fixed, before randomization. Following angrist1996, we classify people into types based on their potential outcomes in intervention and control: “always takers” take up the desired health behavior in intervention and control (effect=0); “compliers” take up in intervention but not control (effect=1); “defiers” do not take up in intervention but take up in control (effect=-1); and “never takers” do not take up in intervention or control (effect=0). We specify the joint distribution of potential outcomes with the numbers of always takers, compliers, defiers, and never takers in the sample. The joint distribution of potential outcomes provides more information than the average effect; it provides the full distribution of effects and disambiguates between zero effects from always and never takers.
@@
What is your best estimate of the numbers of always takers, compliers, defiers, and never takers in the experiment you ran with six people? We estimate with our design-based maximum likelihood decision rule that the sample includes four compliers and two defiers, and we develop a novel visualization of potential outcomes to illustrate the logic behind our answer in Figure 1. We represent each of the six people in the sample with a circle. The left half of each circle represents the potential outcome in intervention, and the right half represents the potential outcome in control. Orange represents “takeup,” white represents “no takeup,” and grey represents “unobserved.”
@@
Rectangles with lines down the middle represent experiments with intervention on the left and control on the right. Inside each experiment, assignment to intervention reveals the potential outcome in intervention—the left of the rectangle reveals the left of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in control is unobserved, so it is grey. Similarly, assignment to control reveals the potential outcome in control—the right of the rectangle reveals the right of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in intervention is unobserved, so it is grey.
@@
The data from your experiment, depicted in the box at the top of Figure 1, show takeup among two of three people in intervention and one of three people in control. Therefore, the estimated marginal distribution of the potential outcome in intervention is 2/3 takeup, and the estimated marginal distribution of the potential outcome in control is 1/3 takeup. These two marginal distributions define the “estimated Fréchet set.” We refer to all joint distributions of potential outcomes with the same marginal distributions as members of the same “Fréchet set.”
@@
The three rows of Figure 1 give the three joint distributions of potential outcomes with the same marginal distributions as the estimated Fréchet set: half circles in which the potential outcomes are unobserved in the data are filled in so that they have the same marginal distributions as the data. Four of the six circles have orange on the left (takeup occurs among 2/3 of people in intervention), and two of the six circles have orange on the right (takeup occurs among 1/3 of people in control). As shown in the three rows, there are three ways to combine the colored half circles into colored full circles, each yielding a different number of defiers (circles with white on the left and orange on the right). This depiction demonstrates how the estimated Fréchet set determines the estimated Fréchet bounds on defiers, which extend here from a lower bound of 0 to an upper bound of 2. How do you decide among these three possible estimates?
Figure 1: In an Experiment with Six People, the Likelihood Varies Among Joint Distributions of Potential Outcomes with the Same Marginal Distributions as the Data (2/3 Takeup in Intervention, 1/3 Takeup in Control) and is Maximized with Four Compliers and Two Defiers
@@
The joint distribution of potential outcomes in the first row is consistent with the imbens1994 assumption of “monotonicity” in the sample, whereby there can be compliers or defiers, but not both. Since the estimated effect of 1/3 is positive, you assume away defiers. You then deduce that the person who takes up in control must be an always taker and the person who does not take up in intervention must be a never taker. Assuming that the randomization guarantees that the count of each type is the same in each arm, the sample includes two always takers and two never takers. The remaining two people must be compliers, so the joint distribution of potential outcomes consistent with the monotonicity assumption includes two always takers, two compliers, and two never takers. However, you are uncomfortable assuming monotonicity because you are concerned that there could be defiers, and you would prefer to decide based on evidence.
@@
The joint distributions of potential outcomes in the next two rows include defiers. The one in the second row includes one always taker, three compliers, one defier, and one never taker. The one in the third row includes four compliers and two defiers. There is a rationale in the literature for deciding on the distribution in the second row—it has one defier, the midpoint (li2019) of the Fréchet bounds on defiers. It can therefore seem surprising that we decide on the distribution in the third row with two defiers.
@@
Each row in Figure 1 provides a rationale for our decision using a “design-based model” that combines the design of the randomization with a model of potential outcomes to specify a data generating process and use it to generate all possible data. As depicted in the top of each row, there are 20 possible randomized assignments of six people that result in three in intervention ($6\text{ choose }3=20$). Each of the six people in the experiment has a separate row. The person’s type and randomized assignment to intervention or control determines the data the experiment reveals, which can be counterfactual to what you observed. The assignments that generate the data you observed—2/3 takeup in intervention and 1/3 takeup in control—are in lighter shading.
@@
The number of randomized assignments that generate the data you observed gives the numerator of the design-based likelihood, reported in the third column. Paraphrasing the baby board book “Statistical Physics for Babies” (ferrie2017), which provides style inspiration for our visualization, physicists refer to the number of ways that you could have seen what you have seen as “entropy.” The design-based likelihood is equal to the entropy—the number of assignments that generate the data (shown in lighter shading)—divided by the number of possible assignments. The first row, with zero defiers, generates the data in eight of 20 possible assignments, so the likelihood is $8/20=40\%$. The second row, with one defier, generates the data in six of 20 possible assignments, so the likelihood is $6/20=30\%$. The last row, with two defiers, generates the data in 12 of 20 possible assignments, so the likelihood is $12/20=60\%$.
@@
Why do some joint distributions of potential outcomes generate the data in more randomized assignments than others and therefore have higher likelihoods than others? The bottom half of each row sorts the six people by type and separately shows how the always takers, compliers, defiers, and never takers must be randomized into intervention and control to generate the data. In the first row, there are three types—always takers, compliers, and never takers—and there must be a balance between intervention and control within each type to generate the data. In the second row, there are four types—always takers, compliers, defiers, and never takers—and there must be an imbalance between intervention and control within each type to generate the data. In the third row, there are only two types—compliers and defiers—and there must be a balance between intervention and control within each type to generate the data.
@@
Likelihoods are higher when the people who generate the data are as similar as possible in two senses: they belong to fewer types, and they have a counterpart of the same type in the opposite arm such that those of the same type are balanced between intervention and control. The people who generate the data are most similar in the third row: they belong to the fewest number of types (only two), and those of the same type are balanced between intervention and control. Comparison of the first and third rows illustrates how having fewer types can increase the likelihood. In the first row, there are two ways to randomize one of two always takers into intervention (2 $2\text{ choose }1$=2), two ways to randomize one of two compliers into intervention (2 $2\text{ choose }1$=2), and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are eight assignments that generate the data ($2\times 2\times 2=8$). In the third row, there are six ways to randomize two of four compliers into intervention ($4\text{ choose }2=6$) and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are 12 assignments that generate the data ($6\times 2=12$). Joint distributions of potential outcomes other than those in Figure 1 can generate the data (the true joint distribution of potential outcomes need not preserve the estimated marginal distributions), but by exhaustive grid search, you determine that the global maximizer of the likelihood is the one depicted in the third row, which only includes two types.
@@
By our proposed maximum likelihood decision rule, you conclude that your best estimate for the joint distribution of potential outcomes in the sample consists of four compliers and two defiers. Using this distribution, you recover the estimated average effect of 1/3—of the six people in the experiment, the four compliers have an effect of positive one, and the two defiers have an effect of negative one ($(4\times 1+2\times-1)/6=1/3$). The design-based model of the experiment enables you to learn about not just the average effect of your intervention, but also about something more primitive: the joint distribution of potential outcomes, which yields the entire distribution of effects.
@@ 2 A Design-Based Model of an Experiment
In this section, we formalize the design-based model of an experiment presented in our stylized example. A sample consists of $n$ subjects, each of whom has a binary potential outcome in intervention $Y_{I}\in\{0,1\}$ and a binary potential outcome in control $Y_{C}\in\{0,1\}$, where $1$ represents takeup and 0 represents no takeup. A subject’s observed outcome depends only on their own potential outcomes and their inclusion in the intervention or control arm, ruling out network-type effects through a “no interference” (cox1958) or “stable unit treatment value” (rubin1980) assumption.
@@
where $\mathbf{1}_{\{\cdot\}}$ is the indicator function.
Subjects in the sample belong to one of four types or “principal strata,” defined by combinations of potential outcomes (frangakis2002). Let $\theta_{11}$ represent the number of always takers $(Y_{I}=1,Y_{C}=1)$, $\theta_{10}$ the number of compliers $(Y_{I}=1,Y_{C}=0)$, $\theta_{01}$ the number of defiers $(Y_{I}=0,Y_{C}=1)$, and $\theta_{00}$ the number of never takers $(Y_{I}=0,Y_{C}=0)$. The collection of these four counts ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ constitutes the joint distribution of potential outcomes in the sample, which we write as counts rather than shares to ease the following likelihood notation. The experimental data ${\boldsymbol{X}}=(X_{I1},X_{I0},X_{C1},X_{C0})$ consist of the counts of subjects who take up in intervention $X_{I1}$, who do not take up in intervention $X_{I0}$, who take up in control $X_{C1}$, and who do not take up in control $X_{C0}$.
We adopt a “design-based” approach, in which the joint distribution of potential outcomes $\boldsymbol{\theta}$} is fixed, but unknown, and all randomness in the experimental data $\boldsymbol{X}$} comes from the random assignment of subjects to either intervention or control. As shown in Appendix A, randomization implies a distribution for the experimental data $\boldsymbol{X}$} given the joint distribution of potential outcomes $\boldsymbol{\theta}$}, yielding a likelihood expression. For Bernoulli randomization, also known as simple randomization or independent and identically distributed (IID) randomization, with probability $p$ of assignment to intervention, the likelihood is:
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$.
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (2)
This last expression appears in copas1973.
@@
In the previous section, we present a likelihood for the joint distribution of potential outcomes in the sample. Each joint distribution of potential outcomes implies marginal distributions of potential outcomes in the intervention and control arms. The marginal distribution of the potential outcome in intervention is the number of subjects for whom $Y=1$ in intervention:
@@
The marginal distribution of the potential outcome in control is the number of subjects for whom $Y=1$ in control:
@@
Given a pair of marginal distributions of potential outcomes in intervention and control, there are numerous joint distributions consistent with them. boole1854, hoeffding1940, and frechet1957 derive bounds on the possible copulas connecting marginal distributions of random variables into joint distributions. By applying these bounds to a pair of marginal distributions of potential outcomes, we can determine the set of all joint distributions of potential outcomes consistent with the pair of marginal distributions. We refer to the set of joint distributions of potential outcomes consistent with a given pair of marginal distributions as the “Fréchet set” of those marginal distributions, $\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)$. In Figure 1, the three rows comprise the Fréchet set for $\theta_{1\bullet}=4$ and $\theta_{\bullet 1}=2$. We index the distributions in a Fréchet set by their number of defiers, $\theta_{01}$, which can take values in the range
$\max\{0,-(\theta_{1\bullet}-\theta_{\bullet 1})\}\leq\theta_{01}\leq\min\{\theta_{\bullet 1},n-\theta_{1\bullet}\}.$ (3)
These inequalities provide the so-called “Fréchet bounds” on the number of defiers. Within these bounds, each value for the number of defiers implies a unique number of always takers, never takers, and compliers consistent with the given marginals:
@@
$\theta_{00}=n-\theta_{1\bullet}-\theta_{01}\Big\}.$
@@
$\hat{\theta}_{1\bullet}({\boldsymbol{x}})$ $=n\left(\frac{x_{I1}}{x_{I1}+x_{I0}}\right)$
$\hat{\theta}_{\bullet 1}({\boldsymbol{x}})$ $=n\left(\frac{x_{C1}}{x_{C1}+x_{C0}}\right).$ (4)
@@
When we report these estimates, we round them if necessary for comparison to our count-based distributions. In our design-based model of an experiment, these are unbiased estimators of the sample marginal distributions of potential outcomes for a completely randomized experiment (before rounding). In a standard sampling-based model of an experiment, dividing by $n$ yields consistent estimates of the population marginal distributions of potential outcomes.
@@
It is well-known that, in the sampling-based model of an experiment, the takeup counts in intervention and control contain no information about the joint distribution of potential outcomes in the population beyond their marginal distributions. In Appendix B, we recreate this result by showing that the likelihood of the population-level joint distribution of potential outcomes is constant, conditional on the population-level marginal distributions. In other words, the likelihood function is flat across every population-level Fréchet set.
@@
In our design-based model of an experiment, however, the data do contain information about the sample joint distribution of potential outcomes beyond the sample marginal distributions. That is, the likelihood function varies with the number of defiers within sample-level Fréchet sets. Figure 2 plots the likelihood values of each joint distribution of potential outcomes in the Fréchet set formed by the estimated marginal distributions from the second empirical example that we will discuss in Section 3. We index the joint distributions in this Fréchet set by their number of defiers along the horizontal axis. The height of the bar is the likelihood of the joint distribution in the Fréchet set with the specified number of defiers. Within this Fréchet set, the likelihood is highest at the distribution with 21 defiers, the upper Fréchet bound on defiers. As the number of defiers increases, the likelihood initially decreases, and then increases. In all empirical examples we have considered, the likelihood maintains a U-shape within the estimated Fréchet set and all other Fréchet sets, and is maximal at either the minimum or maximum number of defiers. If we normalize likelihoods in this Fréchet set to sum to one, the smallest collection of likelihoods to contain 95% of the mass excludes 9 defiers, indicated by the lighter shading.